# Fabian Mauthe

### v2

@@
Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation: Fabian Mauthe, Luca Steinmann, and Peter Zeiler. Overview and analysis of publicly available degradation data sets for tasks within prognostics and health management. In Eirik Bjorheim Abrahamsen, Terje Aven, Frederic Bouder, Roger Flage, and Marja Yloenen, editors, 35th European Safety and Reliability Conference. (Submitted). Research Publishing, 2025.
@@ 1 Prognostics and Health Management
Central to the efficacy of PHM methods is the acquisition and analysis of degradation data, which encapsulates the evolving health condition of ESs over time [3]. Degradation data serves as a rich source of information, offering invaluable insights into the underlying degradation processes, failure modes, and performance trends of ESs [2, 4]. Through the integration of advanced sensing technologies, data analytics, and machine learning algorithms, researchers and practitioners can harness degradation data to develop prognostic models, optimize maintenance schedules, and make informed operational decisions [1].
In the remainder of this paper, the tasks of diagnostics and prognostics are listed, and an overview from [5] of publicly available degradation data sets is presented. The authors aim to update the overview regularly in the future, as new data sets are continually being published in the dynamic research field of PHM.
@@ 2 Tasks of Diagnostics and Prognostics
Extensive and representative degradation data enables addressing a specific task—prognosis—within PHM, with the subdivision of PHM into distinct tasks. Based on [1], [6], and [4], the tasks of diagnostics and prognostics within PHM can be subdivided as follows:
@@
  • Prognosis: Prediction of the future degradation behavior or the current Remaining Useful Life (RUL).
@@ 3 Data Sets for Diagnostics and Prognostics
Table Appendix: Overview of Publicly Available Data Sets in the Appendix: Overview of Publicly Available Data Sets contains an overview of publicly available data sets for the tasks of diagnostics and prognostics within PHM, sorted alphabetically. This is the extended overview from [4], presented in [5].
@@
The detailed PHM-specific analysis of the data sets can be found in [5].
Appendix: Overview of Publicly Available Data Sets
{xltabular}
|p0.6cm|X|p4.1cm|p2.3cm|p1.0cm| Overview of publicly available data sets
@@
Reference to Data Origin Domain
Application and Signal Task URL
@@
Reference to Data Origin Domain
Application and Signal Task URL
[…]
  Continued on next page
@@
1 4TU - Lifecycle ageing tests on commercial 18650 Li ion cell
Trad, Khiem (2021): Lifecycle ageing tests on commercial 18650 Li ion cell @ 25°C and 45°C. Version 1. 4TU.ResearchData. dataset. https://doi.org/10.4121/13739296.v1 Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
@@
Bruinsma, Sietze; Geertsma, Rinze; Loendersloot, Richard; Tinga, Tiedo (2024): Motor Current and Vibration Monitoring Dataset for various Faults in an E-motor-driven Centrifugal Pump. doi: 10.4121/2b61183ec14f-4131-829b-cc4822c369d0.v3 Process technology
@@
Signals: current, vibration, voltage Diagnosis Link
@@ 3 AIDAR Lab - Air Compressor
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2016). Intelligent Condition Based Monitoring using Acoustic Signals for Air Compressors, IEEE Transactions on Reliability, vol. 65, no. 1, pp. 291-309. Mechatronic system
@@
Signals: acoustic emission Diagnosis Link
@@ 4 AIDAR Lab - Drill Bit
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26. Manufacturing process
@@
Signals: vibration Diagnosis Link
@@ 5 Aramis - Data Challenge ESREL2020PSAM15
Francesco Cannarile (Aramis Srl, Italy), Michele Compare (Aramis Srl, Italy and Politecnico di Milano, Italy), Piero Baraldi (Politecnico di Milano, Italy), Zhe Yang (Politecnico di Milano, Italy), Enrico Zio (Politecnico di Milano, Italy and MINES ParisTech, France). Unknown
@@
Signals: vibration Fault detection Link
@@ 6 Backblaze - Hard Drive Stats
Backblaze, Inc. Drive technology
@@
Signals: temperature Fault detection Link
@@ 7 Bearing Data Center - Fan and Bearing
Case Western Reserve University, USA Mechanical component
@@
Signals: acoustic emission Diagnosis Link
@@ 8 Calce - Battery Data Repository
CALCE Center for Advanced Life Cycle Engineering, University of Maryland, website: https://calce.umd.edu/battery-data Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
@@ 9 GitHub - Predictive Maintenance using PySpark
GitHub User: linya9191 Simulation
@@
Signals: unknown Fault detection Link
@@ 10 GitHub - XJTU-SY Bearing Datasets
Biao Wang, Yaguo Lei, Naipeng Li, Ningbo Li (2020). A Hybrid Prognostics Approach for Estimating Remaining Useful Life of Rolling Element Bearings, IEEE Transactions on Reliability, vol. 69, no. 1, pp. 401-412. DOI: 10.1109/TR.2018.2882682. Mechanical component
@@
Signals: vibration Prognosis Link
@@ 11 Kaggle - Air pressure system failures in Scania trucks
Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm Mechatronic system
@@
Signals: anonymized Fault detection Link
@@ 12 Kaggle - Bearings with Varying Degradation Behaviors
Mauthe, F.; Hagmeyer, S.; Zeiler, P. (2025). Holistic simulation model of the temporal degradation of rolling bearings. In E. B. Abrahamsen, T. Aven, F. Bouder, R. Flage, and M. Yloenen (Eds.), Proceedings of the 35nd European Safety and Reliability Conference (Submitted). Research Publishing. Mechanical component
@@
Signals: vibration Prognosis Link
@@ 13 Kaggle - CNC Mill Tool Wear
System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan Manufacturing process
@@
Signals: acceleration, current, feedrate, position, speed, voltage Fault detection Link
@@ 14 Kaggle - Condition Data with Random Recording Time
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
@@
Signals: flow rate, pressure Prognosis Link
@@ 15 Kaggle - E-coating ultrafiltration maintenance
Tinsley Bridge Limited, Sheffield, United Kingdom Manufacturing process
@@
Signals: flow rate, pressure, temperature Prognosis Link
@@ 16 Kaggle - Genesis Demonstrator
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Robotic
@@
Signals: acceleration, current, force, speed Diagnosis Link
@@ 17 Kaggle - Microsoft Azure Predictive Maintenance
Azure AI Notebooks for Predictive Maintenance Unknown
@@
Signals: pressure, speed, vibration, voltage Prognosis Link

### v3

@@
Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation [6]: Fabian Mauthe, Luca Steinmann, Moritz Neu, and Peter Zeiler. Overview and analysis of publicly available degradation data sets for tasks within prognostics and health management. In Proceedings of the 35th European Safety and Reliability Conference and the 33rd Society for Risk Analysis Europe Conference, pages 945–952, Singapore, 2025. Research Publishing. thanks: Citation [2]: Simon Hagmeyer, Fabian Mauthe, and Peter Zeiler. Creation of publicly available data sets for prognostics and diagnostics addressing data scenarios relevant to industrial applications. International Journal of Prognostics and Health Management, 12(2):1–20, 2021.
@@ 1 Prognostics and Health Management
Central to the efficacy of PHM methods is the acquisition and analysis of degradation data, which encapsulates the evolving health condition of ESs over time [1]. Degradation data serves as a rich source of information, offering invaluable insights into the underlying degradation processes, failure modes, and performance trends of ESs [4, 2]. The availability of appropriate degradation data remains a significant challenge in both industrial practice and academic research [10]. Consequently, publicly available degradation data sets are of considerable value. They facilitate the development and empirical benchmarking of PHM methodologies and serve as a basis for demonstrating the practical applicability of the proposed approaches [5, 11, 7].
Although publicly available degradation data sets are of significant importance, the literature addresses this subject to a limited extent. Overviews often consider a restricted range of sources and platforms or only specific applications, resulting in constrained summaries [8, 9]. Analyses usually emphasize the amount of data or data quality, overlooking specific PHM aspects [3]. As a result, the search process for suitable data sets is often very time-consuming for users of PHM methods [6]. Therefore, the objective of this work is to provide a comprehensive overview of publicly available degradation data sets and a detailed PHM-specific analysis. To do this, the analysis presented in [6] is used. This PHM-specific analysis covers aspects that are pertinent to both users of PHM methods and of corresponding data sets. These aspects comprise the application represented in the data set, the originating domain of this application, the specific task the data set is designed to address, and the types of signals contained in the data set [6].
In the remainder of this work, the tasks of diagnostics and prognostics are listed in Section 2, and an overview of publicly available degradation data sets along with a PHM-specific analysis is presented in Section 3. Section 4 concludes the results of this work. The authors aim to update the overview as well as the analysis regularly in the future, as new data sets are continually being published in the dynamic research field of PHM.
@@ 2 Tasks of Diagnostics and Prognostics
Extensive and representative degradation data address a specific task of diagnostics and prognostics within PHM. Based on [10], [9], and [2], these tasks can be subdivided as follows:
@@
  • Prognosis: Prediction of future degradation behavior or the current Remaining Useful Life (RUL).
This list of tasks can also be interpreted as a sequential framework for implementing a comprehensive PHM application, starting with fault detection and ending with prognosis [10]. This can be illustrated by the following example of rolling bearing failure: Fault detection identifies an abnormal vibration pattern. Diagnosis then determines the root cause, for instance, a specific bearing component. The health assessment quantifies how this fault affects the bearing’s condition, while the prognosis predicts the bearing’s RUL. [6]
@@ 3 Data Sets for Diagnostics and Prognostics
3.1 Overview
Due to regular data challenges and several research activities of the PHM community in recent years, new data sets are continuously being published. Therefore, the overview by Mauthe et al. [6] (which is already based on the overview by Hagmeyer et al. [2]) is updated using a similar procedure, i.e., the identical scope of platforms and sources of degradation data sets is considered [2]. As a result, the overview is extended to 110 data sets in total. In accordance with [2] and [6], data sets that focus solely on process quality without including the degradation state are excluded from consideration. Table Appendix A: Overview of Publicly Available Data Sets in Appendix A: Overview of Publicly Available Data Sets contains the extended overview of publicly available data sets for the tasks of diagnostics and prognostics within PHM, sorted alphabetically.
3.2 Data Set Analysis
The detailed PHM-specific analysis of the data sets is based on the procedure presented in [6]. The analysis is intended to shorten the time-consuming search for suitable data sets for its users. In particular, the focus of this analysis is on the following PHM-specific aspects [6]:
  • Application: The application within a data set refers to the object of consideration from which the data originate. Thus, an application can represent a single component, such as a bearing, or an entire system, such as a production line.
  • Domain: The domain from which the application originates. The domains associated with the previous examples would then be, for example, mechanical components or production systems.
  • Task: The task, which is addressed by the corresponding data set. A distinction is made among the tasks described in Section 2: fault detection, diagnosis, health assessment, and prognosis.
  • Signal: The types of signals included in the data set. This refers to signals that offer insight into the health state of the specific application or can be derived from it and used within PHM methods.
Based on these PHM-specific aspects, Mauthe et al. [6] developed a taxonomy that is tailored to the classification of data sets in a PHM-specific manner. This taxonomy categorizes all data sets based on their application and domain of origin, the task addressed by the data set, and the relevant signals contained in the data set. The results using this taxonomy for the 110 publicly available data sets are provided in Table 1 in Appendix B: Classification of Publicly Available Data Sets and in Table 2 in Appendix C: Signals within Publicly Available Data Sets. The respective assignment of the PHM-specific aspects (application, domain, task, and signal) to the individual data sets can also be found in Table Appendix A: Overview of Publicly Available Data Sets in Appendix A: Overview of Publicly Available Data Sets.
Domains and applications:
Applying the presented taxonomy, the 110 publicly available degradation data sets are classified into 11 domains (see Table 1). The mechanical component domain contains the most data sets (26), followed by the electrical component domain (25) and the drive technology domain (12). Apart from these three largest domains, the remaining domains contain between two and eight data sets. Moreover, five data sets have unknown applications. As the respective data sets can still be used to apply PHM methods, they are taken into account and assigned to the unknown domain.
In total, 43 different applications are included in the 11 domains, covering a variety of degradation processes. However, as these applications are spread across 110 data sets, the number of available data sets per application is limited. This is reflected in a more in-depth consideration of Table 1: The applications bearing (17 data sets) and battery (16 data sets) are considered most frequently, originating from the two main domains, mechanical component and electrical component, respectively. The next most frequent applications are aircraft engine, filtration, and production line, each occurring in five data sets. The limited number of data sets per application is further highlighted by the mechatronic system domain, which contains the most different applications (eight), with only one data set per application. Overall, only nine applications are covered by three or more data sets.
Considered tasks within data sets:
The diagnostic and prognostic tasks of the data sets [^1] Certain data sets may be assigned to several tasks. As these tasks build on each other, the highest-ranking task is assigned to a respective data set. Therefore, for each data set, only one task is assigned., as introduced in Section 2, are shown in the third column of Table 1. With 54 assigned data sets, the prognosis task is the most represented, followed by diagnosis with 35 entries, whereas fault detection is assigned to 18 data sets. As in the previous classification, the transition from diagnosis to health assessment, as well as the transition to prognosis, is often not distinct. Accordingly, only three data sets explicitly address the task of health assessment.
For prognosis, the electrical component domain contributes 22 out of 54 data sets, of which 15 address battery applications. The mechanical component and drive technology domains follow with eight and seven prognosis data sets, respectively. Bearing is the second most common application for prognosis (seven data sets), while aircraft engine and filtration each occur in five prognosis data sets. In total, 20 distinct applications are available for prognosis.
For diagnosis, the mechanical component domain contains the most data sets (13 out of 35), while the robotic and mechatronic system domains contain five diagnosis data sets each, and the production system domain contains three. The bearing application is the most frequently used for diagnosis (eight data sets), followed by articulated robot applications (four) and both gear and production line applications (three each). In general, 21 different applications are considered in the 35 data sets for diagnosis, indicating a wide variety of applications.
For fault detection, the mechanical component and production system domains contribute the largest shares, with five and four data sets, respectively. Drive technology and manufacturing process contribute two fault detection data sets each, and the unknown domain contains two fault detection data sets without a specified application. Regarding applications, only bearing, milling, and production line occur in two data sets each; all remaining applications occur only once.
Signals within data sets:
Table 2 summarizes the 22 signals used in at least two of the given data sets. Note that Table 2 only lists signals that are measured directly. Values calculated from these, such as capacity or power, are excluded. Nevertheless, if only calculated signals are included in a data set, they are still listed. Also summarized are signals of a comparable nature, such as speed, including rotation and velocity.
Vibration appears most frequently, occurring in 34 data sets, followed by current and temperature with 33 data sets each, and voltage with 25 data sets. These signals are typically measured within electrical and mechanical component domains (e.g., battery and bearing applications) and therefore are most prevalent in the analyzed data sets. Speed and pressure are also common, appearing in 20 and 18 data sets, respectively. Anonymized signals and flow rate occur in ten data sets each; however, anonymized signals permit specific observations mainly within a data set or application, limiting their utility for general analysis. Regarding the availability of data sets per signal, a limitation similar to applications can be observed: ten of the 22 signals occur in four or fewer data sets.
Across tasks, signal occurrences show characteristic differences. For prognosis, current (22) and temperature (21) are most prevalent, followed by voltage (19) and vibration (14). For diagnosis, vibration dominates (16), followed by speed (11), current (nine), and temperature (eight), indicating that these signals are also relevant beyond purely electrical component data sets.
4 Conclusion
This work updates an overview of publicly available degradation data sets for tasks of diagnostics and prognostics within PHM and structures them using a dedicated taxonomy. The classification makes the current coverage across domains, applications, and tasks explicit and highlights where publicly available data are concentrated versus where only sparse coverage exists. Overall, the results underline that benchmarking is well supported for a few frequently studied use cases and signal modalities, while broader generalization and task-specific evaluations remain limited by uneven coverage. Maintaining and extending this overview will help track how coverage evolves and where additional publicly available data would have the greatest impact.
Appendix A: Overview of Publicly Available Data Sets
@@ 4 Conclusion
Overview of publicly available data sets
@@
Reference to Data Origin
Domain
Application and Signal
Task URL
@@
Reference to Data Origin
Domain
Application and Signal
Task URL
@@
Trad, Khiem (2021): Lifecycle ageing tests on commercial 18650 Li ion cell @ 25°C and 45°C. Version 1. 4TU.ResearchData. dataset. https://doi.org/10.4121/13739296.v1
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
@@
Bruinsma, Sietze; Geertsma, Rinze; Loendersloot, Richard; Tinga, Tiedo (2024): Motor Current and Vibration Monitoring Dataset for various Faults in an E-motor-driven Centrifugal Pump. doi: 10.4121/2b61183ec14f-4131-829b-cc4822c369d0.v3
Process technology
@@
Signals: current, vibration, voltage
Diagnosis Link
@@ 3 AIDAR Lab - Air Compressor
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2016). Intelligent Condition Based Monitoring using Acoustic Signals for Air Compressors, IEEE Transactions on Reliability, vol. 65, no. 1, pp. 291-309.
Mechatronic system
@@
Signals: acoustic emission
Diagnosis Link
@@ 4 AIDAR Lab - Drill Bit
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26.
Manufacturing process
@@
Signals: vibration
Diagnosis Link
@@ 5 Aramis - Data Challenge ESREL2020PSAM15
Francesco Cannarile (Aramis Srl, Italy), Michele Compare (Aramis Srl, Italy and Politecnico di Milano, Italy), Piero Baraldi (Politecnico di Milano, Italy), Zhe Yang (Politecnico di Milano, Italy), Enrico Zio (Politecnico di Milano, Italy and MINES ParisTech, France).
Unknown
@@
Signals: vibration
Fault detection Link
@@ 6 Backblaze - Hard Drive Stats
Backblaze, Inc.
Drive technology
@@
Signals: temperature
Fault detection Link
@@ 7 Bearing Data Center - Fan and Bearing
Case Western Reserve University, USA
Mechanical component
@@
Signals: acoustic emission
Diagnosis Link
@@ 8 Calce - Battery Data Repository
CALCE Center for Advanced Life Cycle Engineering, University of Maryland, website: https://calce.umd.edu/battery-data
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
9 ETH Zurich - Run-to-Failure High-voltage Circuit Breaker Mechanical Test Dataset
Hsu, C.-C. (2024). Run-to-failure high-voltage circuit breaker mechanical test dataset. ETH Zürich Research Collection. DOI: 10.3929/ethz-b-000676480
Electrical component
Circuit breaker (high-voltage)
Signals: current, vibration
Prognosis Link
10 GitHub - Motor current milling machine
Apurv Rajeshkumar Darji (Developer), Dr. Mustafa Demetgül (Academic Supervisor)
Manufacturing process
Milling machine (motor current)
Signals: current
Fault detection Link
@@ 11 GitHub - Predictive Maintenance using PySpark
GitHub User: linya9191
Simulation
@@
Signals: unknown
Fault detection Link
@@ 12 GitHub - XJTU-SY Bearing Datasets
Biao Wang, Yaguo Lei, Naipeng Li, Ningbo Li (2020). A Hybrid Prognostics Approach for Estimating Remaining Useful Life of Rolling Element Bearings, IEEE Transactions on Reliability, vol. 69, no. 1, pp. 401-412. DOI: 10.1109/TR.2018.2882682.
Mechanical component
@@
Signals: vibration
Prognosis Link
13 Harvard Dataverse - Wind Turbine Main Bearing Fatigue Life Prediction
Yucesan, Y. (2019). Wind Turbine Main Bearing Fatigue Life Prediction with PINN [Data set]. Harvard Dataverse. https://doi.org/10.7910/DVN/ENNXLZ
Mechanical component
Bearing
Signals: speed, temperature
Prognosis Link
@@ 14 Kaggle - Air pressure system failures in Scania trucks
Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm
Mechatronic system
@@
Signals: anonymized
Fault detection Link
@@ 15 Kaggle - Bearings with Varying Degradation Behaviors
Mauthe, F.; Hagmeyer, S.; Zeiler, P. (2025). Holistic simulation model of the temporal degradation of rolling bearings. In E. B. Abrahamsen, T. Aven, F. Bouder, R. Flage, and M. Ylönen (Eds.), Proceedings of the 35th European Safety and Reliability Conference and the 33rd Society for Risk Analysis Europe Conference. Research Publishing. DOI: 10.3850/978-981-94-3281-3_ESREL-SRA-E2025-P8028-cd
Mechanical component
@@
Signals: vibration
Prognosis Link
@@ 16 Kaggle - CNC Mill Tool Wear
System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan
Manufacturing process
@@
Signals: acceleration, current, feedrate, position, speed, voltage
Fault detection Link
@@ 17 Kaggle - Condition Data with Random Recording Time
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
@@
Signals: flow rate, pressure
Prognosis Link
@@ 18 Kaggle - E-coating ultrafiltration maintenance
Tinsley Bridge Limited, Sheffield, United Kingdom
Manufacturing process
@@
Signals: flow rate, pressure, temperature
Prognosis Link
@@ 19 Kaggle - Genesis Demonstrator
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Robotic
@@
Signals: acceleration, current, force, speed
Diagnosis Link
@@ 20 Kaggle - Microsoft Azure Predictive Maintenance
Azure AI Notebooks for Predictive Maintenance
Unknown
@@
Signals: pressure, speed, vibration, voltage
Prognosis Link